### v1

# BiMediX2 : Bio-Medical EXpert LMM for Diverse Medical Modalities

Abstract
This paper introduces BiMediX2, a bilingual (Arabic-English) Bio-Medical EXpert Large Multimodal Model (LMM) with a unified architecture that integrates text and visual modalities, enabling advanced image understanding and medical applications. BiMediX2 leverages the Llama3.1 architecture and integrates text and visual capabilities to facilitate seamless interactions in both English and Arabic, supporting text-based inputs and multi-turn conversations involving medical images. The model is trained on an extensive bilingual healthcare dataset consisting of 1.6M samples of diverse medical interactions for both text and image modalities, mixed in Arabic and English. We also propose the first bilingual GPT-4o based medical LMM benchmark named BiMed-MBench. BiMediX2 is benchmarked on both text-based and image-based tasks, achieving state-of-the-art performance across several medical benchmarks. It outperforms recent state-of-the-art models in medical LLM evaluation benchmarks. Our model also sets a new benchmark in multimodal medical evaluations with over 9% improvement in English and over 20% in Arabic evaluations. Additionally, it surpasses GPT-4 by around 9% in UPHILL factual accuracy evaluations and excels in various medical Visual Question Answering, Report Generation, and Report Summarization tasks. The project page, including source code and the trained model, is available at  https://github.com/mbzuai-oryx/BiMediX2.
1 Introduction
Figure 1: Model Performance Comparison on BiMed-MBench: These comparisons are made across different categories, including CT, MRI, CXR, Histology, Gross, and their Arabic counterparts (CT_ar, MRI_ar, CXR_ar, Histology_ar, Gross_ar). The models compared are LLaVA-$p$, LLaVA-Med, BiMediX2, Dragonfly-Med, MiniGPT-Med, and BiomedGPT. Each axis represents the performance score in a specific category, allowing for a visual comparison of how each model performs in bilingual medical contexts.
Recent advancements in healthcare AI, including medical large language models (LLMs) and large multimodal models (LMMs), have shown great potential in improving access to quality medical advice. However, these advancements remain predominantly English-centric, leaving significant gaps in addressing the healthcare needs of non-English-speaking populations. This language bias limits the accessibility and utility of AI-driven medical solutions in regions where languages like Arabic are widely spoken, creating a pressing need for more inclusive and multilingual models. Furthermore, conventional medical LMMs (Zhang et al. (2024); Li et al. (2023); Chen et al. (2024) see Fig 3) generally compromise their advanced medical text-based understanding when integrating multimodal capabilities.
Our project, BiMediX2, strives to address the aforementioned limitations of existing medical LMMs by introducing a Bio-Medical EXpert Large Multimodal Model (LMM) built on the Llama3.1 architecture. It achieves state-of-the-art performance in medical image understanding while also excelling in text-based medical LLM evaluations. BiMediX2 facilitates seamless user interactions in both English and Arabic, leveraging our extensive, purpose-built bilingual multimodal healthcare dataset BiMed-V.
BiMediX2 facilitates medical image understanding along with textual-queries, enabling multi-turn conversations about user-provided medical images. It supports a wide range of imaging modalities, including chest X-rays, CT scans, MRIs, histology slides, and gross pathology images. This multimodal approach ensures comprehensive and accessible medical assistance in a bilingual setting.
Recent advancements in medical LLMs and vision-language models (VLMs) have shown progress in tasks such as multi-turn conversations (MTC), report generation (RG), and specialized domains like radiology (Rad) and pathology (Path) as shown in Tab 1. While models like LLaVa-pp (Rasheed et al. (2024)) and LLaVa-Med (Li et al. (2023)) enable seamless multi-turn conversations, others, such as MiniGPT-Med (Alkhaldi et al. (2024)), are limited to specific modalities like chest X-rays, and BiomedGPT (Zhang et al. (2024)) requires fine-tuned checkpoints for each downstream task. BiMediXv1 (Pieri et al. (2024)) enables multi-turn conversational capabilities and report summarization within a bilingual framework, setting it apart from other medical LLMs. A comprehensive comparison of state-of-the-art models in Clinical LLM evaluations is shown in Fig 3.
Building on this, BiMediX2 expands its functionality to cover a broader spectrum of medical tasks, including Multi-turn Conversations (MTC), Report Summarization (RS), Report Generation (RG), Radiology (Rad), Ophthalmology (Oph), Pathology (Path), Microscopic Analysis (Micro), in a Unified Model (UM) with LLM and VLM capabilities (LLM+VLM). Fig 1 illustrates the performance of state-of-the-art medical LMMs on our BiMed-MBench evaluation in a bilingual context. Its robust bilingual support for Arabic addresses the healthcare needs of Arabic-speaking regions, providing a comprehensive and inclusive solution for diverse, multilingual, and multimodal healthcare applications.
Model MTC RS RG Rad Oph Path Micro LLM+VLM Bil (Ar)
[…]
Model MTC RS RG Rad Oph Path Micro UM LLM+VLM Bil (Ar)
[…]
Table 1: Comparison of Recent Medical LLMs and VLMs. Abbreviations: MTC (Multi-turn conversation), RS (Report Summarization), RG (Report Generation), Rad (Radiology), Oph (Ophthalmology), Path (Pathology), Micro (Microscopic), UM (Unified Model: Single model checkpoint for all downstream tasks), LLM+VLM (Unified LLM + VLM), Bil (Ar) (Bilingual Arabic capabilities).
Contributions
In this work, we present BiMediX2, a bilingual multimodal medical LMM with the following key contributions:
(i) We introduce the first bilingual medical LMM that achieves state-of-the-art results on VLM evaluation benchmarks across various medical image modalities, while also excelling on medical LLM evaluation benchmarks.
(ii) We curated a comprehensive Arabic-English multimodal bilingual instruction set named BiMed-V comprising over 1.6M instructions.
(iii) We introduce the first bilingual GPT-4o-based medical LMM benchmark named BiMed-MBench, consisting of 286 medical queries in English and Arabic across various medical image modalities, fully verified by medical experts.
(iv) Our BiMediX2 LLM outperforms GPT-4 by more than 8% on the USMLE benchmark and by more than 9% in UPHILL factual accuracy evaluations.
(v) Our BiMediX2 LMM achieves state-of-the-art results on BiMed-MBench, with over a 9% improvement in English evaluations and more than a 20% improvement in Arabic evaluations. Furthermore, it excels in medical Visual Question Answering, Report Generation, and Report Summarization tasks.
2 BiMediX2
Figure 2: BiMediX2: Overall Architecture Our model is designed for medical image analysis and bilingual multi-turn conversations. Medical images are processed through a Vision Encoder and aligned with a Projector, while the text inputs are tokenized using the default tokenizer. The resulting tokens are then passed into the language model (Meta Llama 3.1) to generate responses in the prompted language. We only train the language model using LoRA adapters, while the projector is finetuned for medical image-text alignment. A robust data generation framework translates an English data corpus into Arabic using GPT-4o, with verification by a medical expert to ensure accurate and contextually appropriate translations. This approach supports effective training and benchmarking in a bilingual context.
The architecture of BiMediX2 is meticulously designed to facilitate seamless integration of medical image analysis and bilingual multi-turn conversations. At its core, (see Fig 2) the model employs a Vision Encoder (Radford et al. (2021)) to process a diverse array of medical imaging modalities, including chest X-rays, CT scans, MRIs, histology slides, and gross pathology images. This visual data is aligned with textual inputs through a dedicated Projector, ensuring accurate and contextually rich medical image-text mapping. Text inputs are processed using a standard tokenizer, transforming them into the language embedding space of Llama 3.1 (Meta (2024)). This design enables BiMediX2 to generate precise and context-aware responses in either English or Arabic, depending on the user prompt, ensuring flexibility and inclusivity.
Key to BiMediX2’s performance is its modular and efficient training approach. LoRA adapters (Hu et al. (2021)) are utilized to fine-tune the language model while maintaining computational efficiency and minimizing resource demands. The projector is simultaneously fine-tuned to optimize image-text alignment in a medical context. Furthermore, the system is supported by a robust data generation framework, where a comprehensive English data corpus is translated into Arabic using GPT-4o. A random subset of this translation is meticulously verified by bilingual medical experts to ensure clinical relevance and linguistic accuracy. This pipeline enables BiMediX2 to excel in a wide range of medical tasks, including report generation, radiology analysis, pathology insights, and ophthalmological assessments, all within a unified, bilingual, and multimodal framework.
Figure 3: State of the art comparison of models in Clinical LLM Benchmarks
BiMed-V Bilingual and Multimodal Dataset Creation
The BiMed-V dataset is a comprehensive bilingual and multimodal instruction set comprising of 1.6M samples, developed to enhance medical image-text alignment and multimodal understanding. It incorporates a diverse range of publicly available datasets, such as PMC-OA (Lin et al. (2023)), Rad-VQA (Lau et al. (2018)), Path-VQA (He et al. (2020)) and SLAKE (Liu et al. (2021)) complemented by custom-curated data. We also curated 163k VQA samples by repurposing the LLaVA-Med (Li et al. (2023)) 60K-IM dataset, which were tailored to align with real-world medical queries. Furthermore, over 10k samples from the LLaVA-Med pretraining dataset were reformatted into interactive conversations using the Llama 3.1 70B model. A subset of the PMC-OA dataset with short question-answer pairs and multiple-choice questions were added to enhance the dataset’s diversity. Training splits of Rad-VQA, Path-VQA and SLAKE, which typically feature concise answers, were restructured into more detailed responses using the same Llama 3.1 70B model, enhancing the dataset’s depth and usability for complex tasks.
A unique feature of BiMed-V is its bilingual support, facilitated by a multimodal instruction set comprising 326k samples across various medical imaging modalities. This includes 163k Arabic-language samples generated via a comprehensive translation framework (see Fig. 2). English datasets were translated into Arabic using GPT-4o, with rigorous verification of a random subset by bilingual medical experts to ensure clinical relevance and linguistic precision. This hybrid approach balances automation and expert validation, significantly reducing reliance on human medical domain experts while maintaining data quality. Additionally, the inclusion of text-based clinical data from BiMediXv1 (Pieri et al. (2024)) ensures the dataset retains robust language understanding capabilities while expanding its multimodal medical proficiency. This extensive dataset forms the foundation for advanced medical image-text alignment and conversational multimodal applications.
Bilingual and Multimodal Medical Instruction Tuning
To enable BiMediX2’s robust capabilities in both bilingual and multimodal medical tasks, we employ a two-stage training process that ensures precise alignment of visual and textual representations while adapting the language model for complex medical instruction tasks.
Stage 1 Training: Medical Concept Alignment
In the first stage, we finetune the Projector alone to align visual embeddings to the language embedding space. The training utilizes a comprehensive dataset of 467k image-caption pairs sourced from the LLaVA-Med (Li et al. (2023)) pretraining dataset. These pairs span diverse medical imaging modalities and captions that describe clinically relevant features.
Stage 2 Training: Multimodal Medical Instruction Alignment
The second stage finetunes the LoRA(Hu et al. (2021)) adapters within the language model, enhancing its ability to process and generate multimodal medical instructions. For this, we utilize our BiMed-V 1.6M bilingual multimodal instruction set, which comprises meticulously crafted English and Arabic prompts paired with corresponding visual and textual responses. This dataset enables the model to learn nuanced instructions across a wide array of medical domains, from radiology to pathology, in a bilingual context.
Through these two stages, BiMediX2 achieves seamless integration of bilingual and multimodal capabilities, enabling it to deliver accurate and context-aware medical insights in both English and Arabic, tailored to a variety of clinical scenarios.
3 Experiments
In the literature, evaluating medical language models predominantly involves multiple-choice question-answering tasks, with accuracy as the performance metric. We employed the EleutherAI (Gao et al. (2024)) evaluation framework for evaluating text-only medical tasks following (Pieri et al. (2024)). For our bilingual and multimodal evaluations, we utilized several benchmarks and tools. We leveraged our BiMed-MBench for assessing the correctness, helpfulness, relevance and accuracy of the model in a bilingual medical multimodal context. Additionally, we evaluated the model on VQA using datasets such as Path-VQA, SLAKE, and Rad-VQA. For report generation, we used the MIMIC-CXR (Johnson et al. (2019)) dataset, and for report summarization, we used the MIMIC-III (Johnson et al. (2016)) dataset. These evaluations were conducted using the MultiMedEval (Royer et al. (2024)) toolkit, ensuring a comprehensive assessment of the model’s capabilities in both textual and visual medical tasks.
Medical LLM Benchmarks
Model MMLU MedMCQA MedQA USMLE PubmedQA Average
Cli-KG C-Bio C-Med Med-Gen Pro-Med Ana
BioMedGPT-LM-7B 49.4 43.1 41.4 45.0 51.0 45.2 34.8 33.2 31.7 74.0 44.9
BiMediX2 4B 55.1 63.9 47.4 55.0 36.0 52.6 38.1 37.9 47.1 72.2 50.5
LLaVA-Med 59.6 59.7 50.9 59.0 51.5 51.9 44.5 35.7 36.9 74.0 52.4
Dragonfly-Med 65.6 69.4 56.6 69.0 58.4 57.0 49.9 42.8 46.1 75.4 59.0
GPT 3.5 69.8 72.2 61.3 70.0 70.2 56.3 50.1 50.8 49.1 71.6 62.1
Meditron 70B 68.3 77.8 63.6 75.0 74.6 56.3 48.4 53.1 55.4 76.2 64.9
BiMediX2 8B 77.7 79.2 68.8 82.0 74.3 65.9 58.0 57.0 68.6 72.4 70.4
GPT 4 86.0 95.1 76.9 91.0 93.0 80.0 69.5 78.9 83.8 75.2 82.9
Llama3-Med42-70B 84.2 93.1 79.8 91.0 90.1 80.7 72.5 73.8 84.3 80.6 83.0
OpenBioLLM-70B 92.5 93.8 85.6 93.0 93.4 83.7 74.1 68.9 72.0 78.0 83.5
Llama 3.1 70B 83.4 95.1 79.2 93.0 91.5 80.7 71.7 73.8 92.0 77.6 83.8
BiMediX2 70B 86.8 95.1 79.8 94.0 91.5 82.2 70.5 74.3 92.3 79.0 84.6
Table 2: Clinical LLM Evaluation Benchmark
PubMedQA (Jin et al. (2019)) is a question-answering dataset derived from biomedical research papers on PubMed. Given a question sourced from the title of a research paper and a context from the abstract, the task is to provide an answer in the form of ’yes’, ’no’, or ’maybe’, mirroring the conclusion of the paper. Building on prior research, our examination focuses on the PQA-L subset of PubMedQA, consisting of 500 manually annotated QA pairs sourced from PubMed. Answering these questions with accuracy requires thorough reasoning over the biomedical contexts and quantitative data analysis.
MedMCQA (Pal et al. (2022)) is a multiple-choice dataset constructed from questions featured in the Indian AIIMS and NEET PG medical entrance exams. It covers a broad spectrum of medical subjects, evaluating both professional domain knowledge and language comprehension. The dataset includes a test set of 4,183 questions, each with four options.
MedQA (Jin et al. (2021)) is a dataset comprising multiple-choice questions from medical board examinations in the US, Mainland China, and Taiwan. It features two types of questions: concise sentences to assess specific knowledge and extended paragraphs that detail a patient’s condition. Our analysis concentrates on the English portion (USMLE), sourced directly from the National Board of Medical Examiners (NBME), responsible for the assessments required for medical licensure in the U.S., containing 1,273 samples for testing purposes. Engaging with this benchmark involves multiple stages of reasoning and evidence retrieval.
USMLE (Han et al. (2023)) USMLE Self-assessment Test, derived from the United States Medical Licensing Examination (USMLE) self-assessment for Step 1, Step 2, and Step 3, is composed of multiple-choice questions. We employ the version provided by the authors of MedAlpaca, which specifically omits image-based questions and comprises 325 test samples.
The Medical MMLU (Hendrycks et al. (2020)) (Massive Multitask Language Understanding) is a group of six datasets that compiles 1,089 test questions with four options related to different medical subjects. The six MMLU datasets are Clinical Knowledge (Cli-KG), College Biology (C-Bio), College Medicine (C-Med), Medical Genetics (Med-Gen), Professional Medicine (Pro-Med) & Anatomy (Ana).
UPHILL OpenQA (Kaur et al. (2023)) The UPHILL dataset is designed to evaluate the factual accuracy of LLMs in handling health-related queries in English embedded with varying degrees of presuppositions. Our analysis specifically targeted the zero-shot models’ abilities to process and respond to false claims (factually incorrect or debunked by reliable sources) across a spectrum of presupposition levels, a crucial test given the potential real-world implications of misinformation and inaccuracies in the health domain. In this context, the accuracy refers to the model’s effectiveness in accurately refuting false health-related claims at different presupposition levels.
Medical VLM Benchmarks
BiMed-MBench: We introduced the first bilingual GPT-4o-based medical LMM benchmark, consisting of 286 medical queries spanning various medical imaging modalities. The test dataset includes conversational interactions and detailed descriptions for modalities such as chest X-rays (CXR), MRI, histology, gross and CT scans derived from LLaVA-Med (Li et al. (2023)). GPT-4o evaluates the correctness of model responses based on the provided image context and caption. The reference prediction, serving as the upper-bound answer for the teacher model, is generated using GPT-4 (Achiam et al. (2023)). For Arabic evaluations, the ground truth is derived from the upper-bound, translated by GPT-4o, and all test samples are comprehensively verified by medical experts for quality assurance. GPT-4o evaluates the responses by comparing the candidate LMM against the upper-bound for English queries and against the Arabic ground truth for Arabic queries. The evaluation assesses the helpfulness, relevance, accuracy, and level of detail of the responses while maintaining the user language. The scoring is conducted on a scale of 1 to 10, with higher scores indicating better performance. Additionally, GPT-4o provides detailed explanations of the evaluation to offer deeper insights into the models’ performance. Finally, we compute and compare the relative scores of the two candidates.
Visual Question Answering (VQA): To evaluate a model’s performance in Visual Question Answering, three datasets are used: Path-VQA, SLAKE, and Rad-VQA. These datasets include a mix of open-ended and close-ended questions, designed to test the model’s ability to interpret and respond to queries based on medical images. This evaluation ensures the model’s effectiveness in handling a variety of visual and textual inputs.
Report Generation: The MIMIC-CXR (Johnson et al. (2019)) dataset, which includes de-identified radiology reports associated with chest X-rays, is utilized to assess the model’s performance in generating the findings section of medical reports. For this task, the input consists of one or more radiology images related to a case, followed by a prompt such as, “<<<image>>> <<<image>>> Please caption this scan with findings and impressions.” This setup evaluates the model’s ability to generate accurate and coherent medical reports based on visual inputs.
Report Summarization: We conduct an evaluation of report summarization using the MIMIC-III (Johnson et al. (2016)) dataset. The task involves generating the impressions section of a radiology report based on the findings section. To achieve this, the model is provided with the free-text findings along with a task prompt, “Summarize the findings”. This assessment focuses on the model’s ability to distill comprehensive medical information into a concise and precise summary, which is essential for clear and effective medical communication.
4 Results
LLM Evaluation
The performance of BiMediX2 and other models on various medical benchmarks is presented in Fig 3 and Tab 2. Our BiMediX2 70B achieved the highest average score of 84.6%, outperforming other models such as GPT-4 (82.9%) and Llama-3-Med42-70B (83.0%). BiMediX2 70B exhibited superior average performance across multiple datasets, including Medical MMLU, MedMCQA, MedQA, USMLE, and PubMedQA, demonstrating its strong understanding of medical contexts.
Model Conversation Description CXR MRI Histology Gross CT Overall
BiomedGPT 15.3 13.3 16.4 13.0 14.1 14.9 15.8 14.8
LLaVA-pp 34.3 36.6 44.7 33.3 34.7 30.2 31.5 34.9
MiniGPT-Med 37.5 29.6 47.6 32.5 36.3 31.8 29.1 35.4
LLaVA-Med 55.6 43.3 59.5 43.4 54.4 53.9 51.0 52.4
Dragonfly-Med 59.2 34.2 67.0 51.2 53.7 42.6 48.3 52.7
BiMediX2 8B 64.9 54.5 71.7 56.8 62.5 61.4 58.9 62.2
Table 3: BiMed-MBench English Evaluation
Model Conversation Description CXR MRI Histology Gross CT Overall
BiomedGPT 11.1 11.2 11.4 10.8 11.5 11.3 11.1 11.2
MiniGPT-Med 21.6 12.6 23.7 12.7 32.0 15.8 14.9 20.2
LLaVA-Med 23.9 29.4 31.2 25.3 24.8 23.4 26.4 26.2
LLaVA-pp 29.0 27.8 33.2 25.0 33.0 25.8 25.8 28.7
Dragonfly-Med 32.8 19.9 31.9 25.7 33.0 24.0 31.7 29.5
BiMediX2 8B 54.3 36.2 61.4 44.6 51.5 43.5 50.8 50.5
Table 4: BiMed-MBench Arabic Evaluation
UPHILL OpenQA Evaluation
Fig 4 illustrates the performance comparison on the UPHILL OpenQA benchmark. BiMediX2 70B achieved the highest overall factual accuracy of 60.6%, and the second highest being BiMediX2 8B (56.1%), surpassing other models such as GPT-4 (51.5%), Meditron 70B (49.6%), and Med42 (53.5%). This highlights BiMediX2’s effectiveness in discerning and correcting misinformation in medical contexts.
Figure 4: Performance comparison on UPHILL OpenQA (Kaur et al. (2023)), assessing the model’s ability to address false medical claims at different presupposition levels.
BiMed-MBench Evaluation
Tables 3 and 4 present the evaluation results of BiMediX2 and other medical LMMs on the English and Arabic BiMed-MBench benchmark, respectively. BiMediX2 8B demonstrated superior performance in both English (overall score of 62.2%) and Arabic (overall score of 50.5%) evaluations, outperforming other models. Fig 1 compares the performance of state-of-the-art medical LMMs on our BiMed-MBench evaluation in a bilingual context. This indicates BiMediX2’s strong bilingual capabilities and its effectiveness in handling medical conversations and descriptions across different imaging modalities.
Medical VQA Benchmark
Tab 5 shows the performance of BiMediX2 and other models on the Medical VQA benchmark using the MultiMedEval toolkit. Our BiMediX2 8B achieves the highest average score of 0.611, outperforming other models across datasets such as Rad-VQA, Slake-VQA, and Path-VQA. This demonstrates BiMediX2’s proficiency in visual question answering, a critical task in medical diagnostics.
Dataset Metric RadFM LLaVA Med BioMedGPT MiniGPT-Med Phi-3.5 V BiMediX2 4B BiMediX2 8B
Rad-VQA BLEU-1↑ 0.475 0.033 0.044 0.662 0.377 0.501 0.552
closed Q accuracy↑ 0.577 0.545 0.203 0.829 0.618 0.685 0.725
open Q recall↑ 0.407 0.246 0.199 0.546 0.295 0.292 0.363
recall↑ 0.438 0.372 0.199 0.703 0.475 0.511 0.565
open Q accuracy↑ 0.335 0.140 0.150 0.490 0.200 0.225 0.305
F1 ↑ 0.442 0.069 0.064 0.675 0.391 0.516 0.569
Slake-VQA BLEU-1↑ 0.746 0.036 0.175 0.337 0.089 0.625 0.778
closed Q accuracy↑ 0.752 0.512 0.248 0.572 0.535 0.744 0.831
open Q recall↑ 0.758 0.429 0.293 0.308 0.377 0.624 0.763
recall↑ 0.695 0.443 0.260 0.396 0.404 0.664 0.786
open Q accuracy↑ 0.725 0.362 0.259 0.278 0.329 0.567 0.729
F1 ↑ 0.714 0.075 0.192 0.349 0.129 0.641 0.787
Path-VQA BLEU-1↑ 0.257 0.021 0.145 0.296 0.283 0.469 0.587
closed Q accuracy↑ 0.505 0.512 0.260 0.581 0.553 0.708 0.872
open Q recall↑ 0.020 0.116 0.093 0.040 0.063 0.239 0.314
recall↑ 0.221 0.287 0.176 0.311 0.308 0.474 0.593
open Q accuracy↑ 0.005 0.053 0.077 0.019 0.027 0.210 0.282
F1 ↑ 0.232 0.052 0.154 0.299 0.287 0.475 0.595
Average 0.461 0.239 0.177 0.427 0.319 0.509 0.611
Table 5: Medical VQA Benchmark (MultiMedEval Royer et al. (2024))
Report Summarization
Tab 6 presents the report summarization performance on the MIMIC-III dataset. BiMediX2 8B achieved the highest average score of 0.416, surpassing other models like LLaVA-Med (0.398) and Dragonfly-Med (0.110). This average score is derived as a unified metric by re-scaling BLUE-4* and RadCliQ* metrics. This indicates BiMediX2’s effectiveness in generating concise and accurate summaries of medical reports, a vital task for efficient healthcare communication.
Report Generation
Tab 7 shows the report generation performance on the MIMIC-CXR dataset. BiMediX2 8B achieved the highest average score of 0.235, outperforming other models like LLaVA-Med (0.192) and BioMedGPT (0.145). The average score is derived as a unified metric by re-scaling BLUE-4* and RadCliQ* metrics. This highlights BiMediX2’s capability in generating detailed and accurate medical reports from radiology images, a crucial task for diagnostic purposes.
Dataset Metric LLaVA Med Dragonfly-Med BiMediX2 4B BiMediX2 8B
MIMIC-III ROUGE-L↑ 0.185 0.072 0.209 0.205
BLEU-1↑ 0.192 0.062 0.153 0.178
BLEU-4↑* 0.520 0.000 0.410 0.449
F1-RadGraph↑ 0.232 0.000 0.222 0.230
RadCliQ↑* 0.753 0.247 0.923 0.918
CheXbert vector↑ 0.600 0.326 0.633 0.593
METEOR↑ 0.303 0.060 0.264 0.339
Average 0.398 0.110 0.402 0.416
Table 6: Report Summarization (MultiMedEval Royer et al. (2024))
Dataset Metric RadFM LLaVA Med BioMedGPT BiMediX2 4B BiMediX2 8B
MIMIC-CXR Report Generation F1-RadGraph↑ 0.042 0.048 0.000 0.083 0.098
BLEU-1↑ 0.006 0.163 0.003 0.046 0.155
BLEU-4↑* 0.000 0.060 0.000 0.042 0.016
ROUGE-L↑ 0.065 0.125 0.012 0.131 0.153
RadCliQ↑* 0.655 0.660 0.827 0.865 0.860
CheXbert vector↑ 0.197 0.150 0.153 0.205 0.189
METEOR↑ 0.053 0.137 0.016 0.107 0.174
Average 0.145 0.192 0.145 0.211 0.235
Table 7: Report Generation (MultiMedEval Royer et al. (2024))
Qualitative Examples
Medical Image Understanding in a Conversational Context: Fig 5 illustrates the capabilities of the BiMediX2 framework in analyzing medical images and providing detailed, context-aware responses. The top section showcases a conversation involving a sagittal CT scan of the lumbar spine. The model identifies the scan type and explains that it is a vertical slice of the body focusing on the lumbar spine. When asked about abnormalities, the model accurately identifies a fracture in the L4 vertebra, providing an explanation of the condition and its potential causes, such as trauma or stress. The bottom section features a color Doppler ultrasound scan of the left ovary, identifying a potential abnormality a left ovarian cyst with a solid component. Our model explains the imaging technique, names the organ, and discusses the detected abnormality, emphasizing the need for further evaluation. These examples demonstrate BiMediX2’s ability to interpret complex medical images and provide informative responses to assist in clinical decision-making.
Figure 5: Qualitative Examples of our BiMediX2 for Medical Image Understanding in a Conversational Context.
Medical Image Understanding in a Conversational and Bilingual (Arabic) Setting: Fig 6 illustrates BiMediX2’s ability to understand medical imagery and converse in Arabic, showcasing its bilingual capabilities. In the top section, our model accurately identifies the organ and the type of scan, providing clear and precise information that can assist medical professionals in diagnosing and treating liver-related conditions. The middle section shows that our model identifies the body part and the type of scan. And in the bottom section, BiMediX2 accurately identifies the organ, explains the imaging technique, and detects abnormalities, providing valuable insights that can aid in the diagnosis and treatment of conditions related to the female reproductive system. This particular example showcases BiMediX2’s capability to converse in both English and Arabic simultaneously.
Figure 6: Qualitative Examples of our BiMediX2 for Medical Image Understanding in a Conversational and Bilingual (Arabic) Setting: The top section translates to; User: What is the organ shown in the image? BiMediX2: The image shows the liver, which is a large and vital organ in the human body. User: What type of scan is this? BiMediX2: This is a computed tomography (CT) scan of the abdomen. The middle section translates to; User: What type of imaging technique is used in this image? BiMediX2: The imaging technique used in this image is magnetic resonance imaging (MRI). User: What part of the body is shown in the image? BiMediX2: The image shows the head and neck area, with a particular focus on the pituitary gland. And the bottom section translates to; User: What is the organ shown in the image? BiMediX2: The organ shown in the image is the ovary, which is part of the female reproductive system. User: What type of scan is this? BiMediX2: This is an ultrasound image, which is a non-invasive imaging technique that uses high-frequency sound waves to create detailed images of the body’s internal structures. User: What is the abnormality in this ultrasound image? BiMediX2: The abnormality in this ultrasound image is the presence of a multilocular ovarian cyst.
Medical Image Understanding of our BiMediX2 in Multidomain: Fig 7 showcases BiMediX2’s versatility and accuracy across various medical imaging modalities. In the top section, it identifies key structures in a scanning electron micrograph of a mosquito head. In subsequent sections, it correctly interprets an MRI of a parotid tumor, a histology slide of adipose tissue (including stain type), a chest X-ray with pneumothorax, and a CT scan of the abdomen, pinpointing adrenal abnormalities. These examples demonstrate BiMediX2’s capability to analyze and diagnose diverse imaging types, making it a valuable tool across medical specialties.
Figure 7: Qualitative Examples for Medical Image Understanding of our BiMediX2 in Multidomain: Capability of BiMediX2 to understand and interpret various medical imaging modalities and provide accurate responses. The examples include describing the anatomy of a mosquito’s head in a scanning electron micrograph, detecting a left-sided parotid tumor in an MRI scan, recognizing mature adipose tissue in a histology slide, identifying a right-sided pneumothorax in a chest X-ray, and identifying bilateral adrenal hyperplasia in a CT scan of the abdomen. These examples highlight BiMediX2’s versatility and effectiveness in medical image analysis and diagnosis.
5 Conclusion
BiMediX2 represents a leap forward in bilingual, multimodal medical AI, addressing the global need for accessible and inclusive healthcare solutions in both English and Arabic. By integrating text and visual modalities within a unified architecture, it enables seamless multi-turn interactions for diverse medical tasks, including medical image analysis and complex medical conversations. Key contributions include the comprehensive bilingual dataset, BiMed-V, which provides diverse multimodal medical instructions tailored for both languages, and the introduction of BiMed-MBench, the first bilingual GPT-4o-based medical benchmark, which showcases the model’s ability to excel in a wide range of expert-verified medical scenarios. BiMediX2 paves the way for inclusive, multilingual, and multimodal healthcare applications, significantly enhancing the accessibility and quality of medical assistance worldwide.
6 Limitations
Despite its overall improvement, BiMediX2, like other generative language models, may experience issues such as hallucinations, toxicity, and stereotypes. These issues stem from both the inherited limitations of the base models and the nature of the pretraining data. While we have conducted automatic and qualitative evaluations with medical professionals, we acknowledge that our model’s medical diagnoses and recommendations may not always be accurate. Extensive human evaluation is more reliable but expensive and time-consuming. The exploration of alternative solutions remains an important focus for ongoing research.
Currently, our models lack explicit mechanisms to curb undesirable behaviors. Future research will focus on enhancing security and alignment strategies. On a brighter note, we believe that releasing our model weights could contribute to investigating and mitigating risks tied to application uses.
7 Safety and Ethical Implications
We recognize the significant societal impact of BiMediX2 and emphasize the importance of ethical considerations and transparency. This release is intended for research purposes only and is not ready for clinical or commercial use. Ensuring the model’s accuracy and reliability is crucial, as incorrect medical advice could have serious health consequences. Robust validation and quality control measures are essential to minimize errors.
Ethical considerations include protecting patient privacy and ensuring the confidentiality of medical data. The model must comply with relevant data protection regulations and ethical guidelines. Addressing potential biases in the model’s outputs is also critical for ensuring fair and equitable healthcare outcomes.
Collaboration with patients, medical professionals, and ethicists is essential for ethical oversight and further research to ensure safety and accuracy in clinical settings. By acknowledging and addressing these considerations, we can continue to refine BiMediX2 for safe and effective use in healthcare.
8 Acknowledgement
We would like to thank Dr. Omair Mohammed, Dr. Mohammed Zidan and Dr. Vishal Thomas Oommen for their contribution in verification of medical responses.
The computations were enabled by resources provided by LUMI hosted by CSC (Finland) and LUMI consortium, and by Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the NSC.
We are grateful to the Meta Llama Impact Innovation Awards for recognizing BiMediX2, as one of the winners awarded in October 2024. This recognition highlights our commitment to advancing AI-driven healthcare solutions.

### v2

@@
 Abstract
-This paper introduces BiMediX2, a bilingual (Arabic-English) Bio-Medical EXpert Large Multimodal Model (LMM) with a unified architecture that integrates text and visual modalities, enabling advanced image understanding and medical applications.
-BiMediX2 leverages the Llama3.1 architecture and integrates text and visual capabilities to facilitate seamless interactions in both English and Arabic, supporting text-based inputs and multi-turn conversations involving medical images.
-The model is trained on an extensive bilingual healthcare dataset consisting of 1.6M samples of diverse medical interactions for both text and image modalities, mixed in Arabic and English.
-We also propose the first bilingual GPT-4o based medical LMM benchmark named BiMed-MBench.
-BiMediX2 is benchmarked on both text-based and image-based tasks, achieving state-of-the-art performance across several medical benchmarks.
-It outperforms recent state-of-the-art models in medical LLM evaluation benchmarks.
-Our model also sets a new benchmark in multimodal medical evaluations with over 9% improvement in English and over 20% in Arabic evaluations.
-Additionally, it surpasses GPT-4 by around 9% in UPHILL factual accuracy evaluations and excels in various medical Visual Question Answering, Report Generation, and Report Summarization tasks.
-The project page, including source code and the trained model, is available at  https://github.com/mbzuai-oryx/BiMediX2.
+We introduce BiMediX2, a bilingual (Arabic-English) Bio-Medical EXpert Large Multimodal Model that supports text-based and image-based medical interactions.
+It enables multi-turn conversation in Arabic and English and supports diverse medical imaging modalities, including radiology, CT, and histology.
+To train BiMediX2, we curate BiMed-V, an extensive Arabic-English bilingual healthcare dataset consisting of 1.6M samples of diverse medical interactions.
+This dataset supports a range of medical Large Language Model (LLM) and Large Multimodal Model (LMM) tasks, including multi-turn medical conversations, report generation, and visual question answering (VQA).
+We also introduce BiMed-MBench, the first Arabic-English medical LMM evaluation benchmark, verified by medical experts.
+BiMediX2 demonstrates excellent performance across multiple medical LLM and LMM benchmarks, achieving state-of-the-art results compared to other open-sourced models.
+On BiMed-MBench, BiMediX2 outperforms existing methods by over 9% in English and more than 20% in Arabic evaluations.
+Additionally, it surpasses GPT-4 by approximately 9% in UPHILL factual accuracy evaluations and excels in various medical VQA, report generation, and report summarization tasks.
+Our trained models, instruction set, and source code are available at – https://github.com/mbzuai-oryx/BiMediX2
+BiMediX2: Bio-Medical EXpert LMM for Diverse Medical Modalities
+Sahal Shaji Mullappilly1thanks: Equal Contribution,  Mohammed Irfan Kurpath [^1] footnotemark: 1,  Sara Pieri1  Saeed Yahya Alseiari5,  Shanavas Cholakkal6,  Khaled Aldahmani3,4,  Fahad Khan1,2  Rao Anwer1,  Salman Khan1, Timothy Baldwin1,  Hisham Cholakkal1 1Mohamed Bin Zayed University of Artificial Intelligence (MBZUAI), 2Linköping University 3Shaikh Tahnoon bin Mohammed Medical City (STMC), 4Tawam Hospital 5Sheikh Shakhbout Medical City (SSMC), 6Govt Medical College Kozhikode
 1 Introduction
-Figure 1: Model Performance Comparison on BiMed-MBench: These comparisons are made across different categories, including CT, MRI, CXR, Histology, Gross, and their Arabic counterparts (CT_ar, MRI_ar, CXR_ar, Histology_ar, Gross_ar).
-The models compared are LLaVA-$p$, LLaVA-Med, BiMediX2, Dragonfly-Med, MiniGPT-Med, and BiomedGPT.
-Each axis represents the performance score in a specific category, allowing for a visual comparison of how each model performs in bilingual medical contexts.
-Recent advancements in healthcare AI, including medical large language models (LLMs) and large multimodal models (LMMs), have shown great potential in improving access to quality medical advice.
-However, these advancements remain predominantly English-centric, leaving significant gaps in addressing the healthcare needs of non-English-speaking populations.
-This language bias limits the accessibility and utility of AI-driven medical solutions in regions where languages like Arabic are widely spoken, creating a pressing need for more inclusive and multilingual models.
-Furthermore, conventional medical LMMs (Zhang et al.
-(2024); Li et al.
-(2023); Chen et al.
-(2024) see Fig 3) generally compromise their advanced medical text-based understanding when integrating multimodal capabilities.
-Our project, BiMediX2, strives to address the aforementioned limitations of existing medical LMMs by introducing a Bio-Medical EXpert Large Multimodal Model (LMM) built on the Llama3.1 architecture.
-It achieves state-of-the-art performance in medical image understanding while also excelling in text-based medical LLM evaluations.
-BiMediX2 facilitates seamless user interactions in both English and Arabic, leveraging our extensive, purpose-built bilingual multimodal healthcare dataset BiMed-V.
-BiMediX2 facilitates medical image understanding along with textual-queries, enabling multi-turn conversations about user-provided medical images.
-It supports a wide range of imaging modalities, including chest X-rays, CT scans, MRIs, histology slides, and gross pathology images.
-This multimodal approach ensures comprehensive and accessible medical assistance in a bilingual setting.
-Recent advancements in medical LLMs and vision-language models (VLMs) have shown progress in tasks such as multi-turn conversations (MTC), report generation (RG), and specialized domains like radiology (Rad) and pathology (Path) as shown in Tab 1.
-While models like LLaVa-pp (Rasheed et al.
-(2024)) and LLaVa-Med (Li et al.
-(2023)) enable seamless multi-turn conversations, others, such as MiniGPT-Med (Alkhaldi et al.
-(2024)), are limited to specific modalities like chest X-rays, and BiomedGPT (Zhang et al.
-(2024)) requires fine-tuned checkpoints for each downstream task.
-BiMediXv1 (Pieri et al.
-(2024)) enables multi-turn conversational capabilities and report summarization within a bilingual framework, setting it apart from other medical LLMs.
-A comprehensive comparison of state-of-the-art models in Clinical LLM evaluations is shown in Fig 3.
-Building on this, BiMediX2 expands its functionality to cover a broader spectrum of medical tasks, including Multi-turn Conversations (MTC), Report Summarization (RS), Report Generation (RG), Radiology (Rad), Ophthalmology (Oph), Pathology (Path), Microscopic Analysis (Micro), in a Unified Model (UM) with LLM and VLM capabilities (LLM+VLM).
-Fig 1 illustrates the performance of state-of-the-art medical LMMs on our BiMed-MBench evaluation in a bilingual context.
-Its robust bilingual support for Arabic addresses the healthcare needs of Arabic-speaking regions, providing a comprehensive and inclusive solution for diverse, multilingual, and multimodal healthcare applications.
+Figure 1: Performance comparison on BiMed-MBench.
+The comparison is conducted across different tasks and modalities, including CT, MRI, CXR, Histology, and Gross, along with their Arabic counterparts (CT_ar, MRI_ar, CXR_ar, Histology_ar, and Gross_ar).
+Each axis represents the performance score for a specific category, highlighting BiMediX2’s superior performance across diverse tasks and modalities in both English and Arabic.
 Model MTC RS RG Rad Oph Path Micro LLM+VLM Bil (Ar)
+Meditron Chen et al. (2023)
+Med42 Christophe et al. (2024)
+OpenBioLLM Ankit Pal (2024)
+Llama3.1 Meta (2024)
+BiMediXv1 Pieri et al. (2024)
+Model MTC RS RG Rad Oph Path Micro UM LLM+VLM Bil (Ar)
+LLaVA-pp Rasheed et al. (2024)
+MiniGPT-Med Alkhaldi et al. (2024)
+MAIRA-2 Bannur et al. (2024)
+BioMedGPT Zhang et al. (2024)
+LLaVA-Med Li et al. (2023)
+Dragonfly-Med Chen et al. (2024)
 […]
-[…]
-[…]
-[…]
-[…]
-Model MTC RS RG Rad Oph Path Micro UM LLM+VLM Bil (Ar)
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 1: Comparison of Recent Medical LLMs and VLMs.
+Table 1: Comparison of tasks and modalities addressed by recent medical LLMs and VLMs.
 Abbreviations: MTC (Multi-turn conversation), RS (Report Summarization), RG (Report Generation), Rad (Radiology), Oph (Ophthalmology), Path (Pathology), Micro (Microscopic), UM (Unified Model: Single model checkpoint for all downstream tasks), LLM+VLM (Unified LLM + VLM), Bil (Ar) (Bilingual Arabic capabilities).
-Contributions
-In this work, we present BiMediX2, a bilingual multimodal medical LMM with the following key contributions:
-(i) We introduce the first bilingual medical LMM that achieves state-of-the-art results on VLM evaluation benchmarks across various medical image modalities, while also excelling on medical LLM evaluation benchmarks.
-(ii) We curated a comprehensive Arabic-English multimodal bilingual instruction set named BiMed-V comprising over 1.6M instructions.
-(iii) We introduce the first bilingual GPT-4o-based medical LMM benchmark named BiMed-MBench, consisting of 286 medical queries in English and Arabic across various medical image modalities, fully verified by medical experts.
-(iv) Our BiMediX2 LLM outperforms GPT-4 by more than 8% on the USMLE benchmark and by more than 9% in UPHILL factual accuracy evaluations.
-(v) Our BiMediX2 LMM achieves state-of-the-art results on BiMed-MBench, with over a 9% improvement in English evaluations and more than a 20% improvement in Arabic evaluations.
-Furthermore, it excels in medical Visual Question Answering, Report Generation, and Report Summarization tasks.
-2 BiMediX2
+Recently, medical Large Language Models (LLMs) and medical Large Multimodal Models (LMMs) have shown promising results as conversational assistants for improving accessibility to quality medical advice.
+However, most medical Vision-Language Models (VLMs) referred to interchangeably as medical LMMs in this paper often compromise their text-based understanding (i.e., medical LLM performance) when integrating multimodal capabilities, making it challenging to interact seamlessly with users (see LLM+VLM column in Tab.
+1).
+This is particularly challenging when users initially ask general medical queries in text format and later follow up with questions related to user-provided medical images introduced mid-conversation to provide additional supporting information.
+In addition to supporting diverse LLM and LMM tasks, such as multi-turn conversations and report generation, it is also desirable to support various medical image modalities, such as radiology and pathology, in a unified model.
+However, as shown in Tab.
+1, state-of-the-art medical LMMs, such as MiniGPT-Med Alkhaldi et al.
+(2024), are either restricted to a limited set of medical image modalities, such as radiology, or require separately fine-tuned models for each downstream task (e.g., BiomedGPT Zhang et al.
+(2024)), posing significant challenges for real-world deployment.
+Moreover, advancements in medical LLMs and LMMs remain predominantly English-centric, leaving significant gaps for non-English-speaking populations, particularly in languages like Arabic, spoken by over 400 million people.
+Unlike most state-of-the-art medical LLMs, BiMediX Pieri et al.
+(2024) (referred to as BiMediXv1 in this paper for clarity) supports Arabic-English bilingual interactions and offers diverse text-based interaction capabilities, such as multi-turn conversations, report summarization, and question-answering.
+However, it is limited to text modality and lacks medical image understanding capabilities.
+To address the aforementioned limitations of existing medical LMMs, we propose BiMediX2, a Bio-Medical EXpert Large Multimodal Model that supports diverse medical tasks and modalities while also facilitating seamless user interactions in both English and Arabic.
+1.1 Contributions
 Figure 2: BiMediX2: Overall Architecture Our model is designed for medical image analysis and bilingual multi-turn conversations.
@@
 We only train the language model using LoRA adapters, while the projector is finetuned for medical image-text alignment.
-A robust data generation framework translates an English data corpus into Arabic using GPT-4o, with verification by a medical expert to ensure accurate and contextually appropriate translations.
-This approach supports effective training and benchmarking in a bilingual context.
-The architecture of BiMediX2 is meticulously designed to facilitate seamless integration of medical image analysis and bilingual multi-turn conversations.
-At its core, (see Fig 2) the model employs a Vision Encoder (Radford et al.
-(2021)) to process a diverse array of medical imaging modalities, including chest X-rays, CT scans, MRIs, histology slides, and gross pathology images.
-This visual data is aligned with textual inputs through a dedicated Projector, ensuring accurate and contextually rich medical image-text mapping.
-Text inputs are processed using a standard tokenizer, transforming them into the language embedding space of Llama 3.1 (Meta (2024)).
-This design enables BiMediX2 to generate precise and context-aware responses in either English or Arabic, depending on the user prompt, ensuring flexibility and inclusivity.
-Key to BiMediX2’s performance is its modular and efficient training approach.
-LoRA adapters (Hu et al.
-(2021)) are utilized to fine-tune the language model while maintaining computational efficiency and minimizing resource demands.
-The projector is simultaneously fine-tuned to optimize image-text alignment in a medical context.
-Furthermore, the system is supported by a robust data generation framework, where a comprehensive English data corpus is translated into Arabic using GPT-4o.
-A random subset of this translation is meticulously verified by bilingual medical experts to ensure clinical relevance and linguistic accuracy.
-This pipeline enables BiMediX2 to excel in a wide range of medical tasks, including report generation, radiology analysis, pathology insights, and ophthalmological assessments, all within a unified, bilingual, and multimodal framework.
-Figure 3: State of the art comparison of models in Clinical LLM Benchmarks
-BiMed-V Bilingual and Multimodal Dataset Creation
-The BiMed-V dataset is a comprehensive bilingual and multimodal instruction set comprising of 1.6M samples, developed to enhance medical image-text alignment and multimodal understanding.
-It incorporates a diverse range of publicly available datasets, such as PMC-OA (Lin et al.
-(2023)), Rad-VQA (Lau et al.
-(2018)), Path-VQA (He et al.
-(2020)) and SLAKE (Liu et al.
-(2021)) complemented by custom-curated data.
-We also curated 163k VQA samples by repurposing the LLaVA-Med (Li et al.
-(2023)) 60K-IM dataset, which were tailored to align with real-world medical queries.
-Furthermore, over 10k samples from the LLaVA-Med pretraining dataset were reformatted into interactive conversations using the Llama 3.1 70B model.
-A subset of the PMC-OA dataset with short question-answer pairs and multiple-choice questions were added to enhance the dataset’s diversity.
-Training splits of Rad-VQA, Path-VQA and SLAKE, which typically feature concise answers, were restructured into more detailed responses using the same Llama 3.1 70B model, enhancing the dataset’s depth and usability for complex tasks.
-A unique feature of BiMed-V is its bilingual support, facilitated by a multimodal instruction set comprising 326k samples across various medical imaging modalities.
-This includes 163k Arabic-language samples generated via a comprehensive translation framework (see Fig.
-2).
-English datasets were translated into Arabic using GPT-4o, with rigorous verification of a random subset by bilingual medical experts to ensure clinical relevance and linguistic precision.
-This hybrid approach balances automation and expert validation, significantly reducing reliance on human medical domain experts while maintaining data quality.
-Additionally, the inclusion of text-based clinical data from BiMediXv1 (Pieri et al.
-(2024)) ensures the dataset retains robust language understanding capabilities while expanding its multimodal medical proficiency.
-This extensive dataset forms the foundation for advanced medical image-text alignment and conversational multimodal applications.
-Bilingual and Multimodal Medical Instruction Tuning
-To enable BiMediX2’s robust capabilities in both bilingual and multimodal medical tasks, we employ a two-stage training process that ensures precise alignment of visual and textual representations while adapting the language model for complex medical instruction tasks.
-Stage 1 Training: Medical Concept Alignment
-In the first stage, we finetune the Projector alone to align visual embeddings to the language embedding space.
-The training utilizes a comprehensive dataset of 467k image-caption pairs sourced from the LLaVA-Med (Li et al.
-(2023)) pretraining dataset.
-These pairs span diverse medical imaging modalities and captions that describe clinically relevant features.
-Stage 2 Training: Multimodal Medical Instruction Alignment
-The second stage finetunes the LoRA(Hu et al.
-(2021)) adapters within the language model, enhancing its ability to process and generate multimodal medical instructions.
-For this, we utilize our BiMed-V 1.6M bilingual multimodal instruction set, which comprises meticulously crafted English and Arabic prompts paired with corresponding visual and textual responses.
-This dataset enables the model to learn nuanced instructions across a wide array of medical domains, from radiology to pathology, in a bilingual context.
-Through these two stages, BiMediX2 achieves seamless integration of bilingual and multimodal capabilities, enabling it to deliver accurate and context-aware medical insights in both English and Arabic, tailored to a variety of clinical scenarios.
-3 Experiments
-In the literature, evaluating medical language models predominantly involves multiple-choice question-answering tasks, with accuracy as the performance metric.
-We employed the EleutherAI (Gao et al.
-(2024)) evaluation framework for evaluating text-only medical tasks following (Pieri et al.
-(2024)).
-For our bilingual and multimodal evaluations, we utilized several benchmarks and tools.
-We leveraged our BiMed-MBench for assessing the correctness, helpfulness, relevance and accuracy of the model in a bilingual medical multimodal context.
-Additionally, we evaluated the model on VQA using datasets such as Path-VQA, SLAKE, and Rad-VQA.
-For report generation, we used the MIMIC-CXR (Johnson et al.
-(2019)) dataset, and for report summarization, we used the MIMIC-III (Johnson et al.
-(2016)) dataset.
-These evaluations were conducted using the MultiMedEval (Royer et al.
-(2024)) toolkit, ensuring a comprehensive assessment of the model’s capabilities in both textual and visual medical tasks.
-Medical LLM Benchmarks
+BiMediX2 follows a two-stage training pipeline.
+Stage-1 aligns medical visual concepts using 467K image-caption pairs.
+Stage-2 performs multimodal medical instruction tuning with our proposed BiMed-V 1.6M bilingual instructions comprising both image-text and text-only medical instructions.
+We introduce BiMediX2 an Arabic-English bilingual medical LMM that supports broader spectrum of text-based and multimodal medical tasks, including Multi-Turn Conversations (MTC), Report Summarization (RS), Report Generation (RG), Radiology (Rad), Ophthalmology (Oph), Pathology (Path), Microscopic Analysis (Micro), in a Unified Model (UM) with LLM and VLM capabilities (LLM+VLM) as shown in Tab.
+1.
+It supports a wide range of imaging modalities, including Chest X-rays, CT scans, MRIs, Histology slides, and Gross pathology images (see Fig.
+1).
+The key contributions of our work can be summarized as: (i) We curated a comprehensive Arabic-English bilingual healthcare specialized instruction set named BiMed-V comprising over 1.6M instructions including text-only and image-text instructions across diverse medical image modalities.
+(ii) We introduce the first Arabic-English bilingual medical LMM evaluation benchmark named BiMed-MBench, consisting of 386 medical queries in English and Arabic across various medical image modalities, fully verified by medical experts.
+(iii) We instruction-tune Llama3.1 LLM with our BiMed-V to obtain the first Arabic-English bilingual medical LMM achieving state-of-the-art results among open-sourced models on various VLM evaluation benchmarks across diverse medical image modalities while also excelling on medical LLM evaluation benchmarks in both English and Arabic.
+This is achieved through our multistage training pipeline and carefully curated instruction set, which balances text-only and multimodal medical instructions.
+(iv) We perform comprehensive evaluation of our model across 12 diverse evaluation benchmarks.
+In addition to achieving promising results on our bilingual BiMed-MBench, BiMediX2 achieves excellent results for visual question answering on Path-VQA, SLAKE and Rad-VQA evaluation benchmarks, and report generation task on MIMIC CXR and report summarization on MIMIC-III benchmarks.
+Additionally, BiMediX2 achieves favorable results on diverse medical LLM benchmarks, such Medical MMLU, MedMCQA, MedQA, USMLE, and PubMedQA datasets, and it also shows robustness in discerning and correcting misinformation in medical context on the UPHILL OpenQA Evaluation.
+Fig 1 illustrates the performance of SOTA medical LMMs on our BiMed-MBench evaluation.
+BiMediX2 achieves SOTA results on BiMed-MBench, with over a 9% improvement in English evaluations and more than a 20% improvement in Arabic evaluations.
+Furthermore, it excels in medical Visual Question Answering, Report Generation, and Report Summarization tasks.
+Similarly, BiMediX2 outperforms GPT-4 by more than 8% on the USMLE benchmark and by more than 9% in UPHILL factual accuracy evaluations.
+2 BiMediX2
 Model MMLU MedMCQA MedQA USMLE PubmedQA Average
@@
 BioMedGPT-LM-7B 49.4 43.1 41.4 45.0 51.0 45.2 34.8 33.2 31.7 74.0 44.9
-BiMediX2 4B 55.1 63.9 47.4 55.0 36.0 52.6 38.1 37.9 47.1 72.2 50.5
 LLaVA-Med 59.6 59.7 50.9 59.0 51.5 51.9 44.5 35.7 36.9 74.0 52.4
 Dragonfly-Med 65.6 69.4 56.6 69.0 58.4 57.0 49.9 42.8 46.1 75.4 59.0
+Apollo-7B 64.2 73.6 59.5 70.0 70.6 61.5 54.4 50.2 52.3 39.0 59.5
 GPT 3.5 69.8 72.2 61.3 70.0 70.2 56.3 50.1 50.8 49.1 71.6 62.1
 Meditron 70B 68.3 77.8 63.6 75.0 74.6 56.3 48.4 53.1 55.4 76.2 64.9
-BiMediX2 8B 77.7 79.2 68.8 82.0 74.3 65.9 58.0 57.0 68.6 72.4 70.4
+BiMediXv1 78.9 86.1 68.2 85.0 80.5 74.1 62.7 62.8 66.8 80.2 74.5
+Apollo-72B 82.3 90.3 77.5 85.0 86.0 70.4 66.7 65.3 74.2 78.8 77.6
 GPT 4 86.0 95.1 76.9 91.0 93.0 80.0 69.5 78.9 83.8 75.2 82.9
@@
 Llama 3.1 70B 83.4 95.1 79.2 93.0 91.5 80.7 71.7 73.8 92.0 77.6 83.8
-BiMediX2 70B 86.8 95.1 79.8 94.0 91.5 82.2 70.5 74.3 92.3 79.0 84.6
+[…]
+[…]
+[…]
 Table 2: Clinical LLM Evaluation Benchmark
-PubMedQA (Jin et al.
-(2019)) is a question-answering dataset derived from biomedical research papers on PubMed.
-Given a question sourced from the title of a research paper and a context from the abstract, the task is to provide an answer in the form of ’yes’, ’no’, or ’maybe’, mirroring the conclusion of the paper.
-Building on prior research, our examination focuses on the PQA-L subset of PubMedQA, consisting of 500 manually annotated QA pairs sourced from PubMed.
-Answering these questions with accuracy requires thorough reasoning over the biomedical contexts and quantitative data analysis.
-MedMCQA (Pal et al.
-(2022)) is a multiple-choice dataset constructed from questions featured in the Indian AIIMS and NEET PG medical entrance exams.
-It covers a broad spectrum of medical subjects, evaluating both professional domain knowledge and language comprehension.
-The dataset includes a test set of 4,183 questions, each with four options.
-MedQA (Jin et al.
-(2021)) is a dataset comprising multiple-choice questions from medical board examinations in the US, Mainland China, and Taiwan.
-It features two types of questions: concise sentences to assess specific knowledge and extended paragraphs that detail a patient’s condition.
-Our analysis concentrates on the English portion (USMLE), sourced directly from the National Board of Medical Examiners (NBME), responsible for the assessments required for medical licensure in the U.S., containing 1,273 samples for testing purposes.
-Engaging with this benchmark involves multiple stages of reasoning and evidence retrieval.
-USMLE (Han et al.
-(2023)) USMLE Self-assessment Test, derived from the United States Medical Licensing Examination (USMLE) self-assessment for Step 1, Step 2, and Step 3, is composed of multiple-choice questions.
-We employ the version provided by the authors of MedAlpaca, which specifically omits image-based questions and comprises 325 test samples.
-The Medical MMLU (Hendrycks et al.
-(2020)) (Massive Multitask Language Understanding) is a group of six datasets that compiles 1,089 test questions with four options related to different medical subjects.
-The six MMLU datasets are Clinical Knowledge (Cli-KG), College Biology (C-Bio), College Medicine (C-Med), Medical Genetics (Med-Gen), Professional Medicine (Pro-Med) & Anatomy (Ana).
-UPHILL OpenQA (Kaur et al.
-(2023)) The UPHILL dataset is designed to evaluate the factual accuracy of LLMs in handling health-related queries in English embedded with varying degrees of presuppositions.
-Our analysis specifically targeted the zero-shot models’ abilities to process and respond to false claims (factually incorrect or debunked by reliable sources) across a spectrum of presupposition levels, a crucial test given the potential real-world implications of misinformation and inaccuracies in the health domain.
-In this context, the accuracy refers to the model’s effectiveness in accurately refuting false health-related claims at different presupposition levels.
-Medical VLM Benchmarks
-BiMed-MBench: We introduced the first bilingual GPT-4o-based medical LMM benchmark, consisting of 286 medical queries spanning various medical imaging modalities.
-The test dataset includes conversational interactions and detailed descriptions for modalities such as chest X-rays (CXR), MRI, histology, gross and CT scans derived from LLaVA-Med (Li et al.
-(2023)).
+The architecture of BiMediX2 is designed to facilitate seamless integration of medical image analysis and bilingual multi-turn conversations.
+At its core, (see Fig 2) the model employs a Vision Encoder Radford et al.
+(2021) to process a diverse array of medical imaging modalities, including chest X-rays, CT scans, MRIs, histology slides, and gross pathology images.
+This visual data is aligned with textual inputs through a dedicated Projector, ensuring accurate and contextually rich medical image-text mapping following Liu et al.
+(2023).
+As shown in Fig 2 we use the ‘<<image>>’ token as a place holder to encode the visual features for multimodal medical instructions.
+For text based medical data the inputs are processed using a standard tokenizer, transforming them into the language embedding space of Llama 3.1 Meta (2024).
+This design enables BiMediX2 to generate precise and context-aware responses in either English or Arabic, depending on the user prompt and supports multimodal interactions while preserving the medical LLM capabilities.
+Key to BiMediX2’s performance is its modular and efficient training approach.
+LoRA adapters Hu et al.
+(2021) are utilized to fine-tune the language model while maintaining computational efficiency and minimizing resource demands.
+The projector is simultaneously fine-tuned to optimize image-text alignment in a medical context.
+Furthermore, the system is supported by a robust data generation framework, where a comprehensive English data corpus is translated into Arabic using GPT-4o.
+A random subset of this translation is meticulously verified by bilingual medical experts to ensure clinical relevance and linguistic accuracy.
+This pipeline enables BiMediX2 to excel in medical tasks, including report generation, radiology analysis, pathology insights, and ophthalmological assessments, in a unified, bilingual, and multimodal framework.
+2.1 BiMed-V: Multimodal, Bilingual Dataset
+The BiMed-V dataset is a comprehensive bilingual and multimodal instruction set comprising of 1.6M samples, developed to enhance medical image-text alignment and multimodal understanding.
+It incorporates a diverse range of publicly available datasets, such as PMC-OA Lin et al.
+(2023), Rad-VQA Lau et al.
+(2018), Path-VQA He et al.
+(2020) and SLAKE Liu et al.
+(2021) complemented by custom-curated data.
+We also curated 163k VQA samples by repurposing the LLaVA-Med Li et al.
+(2023) 60K-IM dataset.
+Furthermore, over 10k samples from the LLaVA-Med pretraining dataset were reformatted into interactive conversations using the Llama 3.1 70B model.
+A subset of the PMC-OA dataset with short question-answer pairs and multiple-choice questions were added to enhance the dataset’s diversity.
+Training splits of Rad-VQA, Path-VQA and SLAKE, which typically feature concise answers, were restructured into more detailed responses using the same Llama 3.1 70B model, enhancing the dataset’s depth and usability for complex tasks.
+A unique feature of BiMed-V is its bilingual support, facilitated by a multimodal open-ended instruction set comprising 326k samples across various medical imaging modalities.
+This includes 163k Arabic-language samples generated via a comprehensive translation framework (see Fig.
+5).
+English datasets were translated into Arabic using GPT-4o, with verification of a random subset by bilingual medical experts to ensure clinical relevance and linguistic precision.
+This hybrid approach balances automation and expert validation, significantly reducing reliance on human medical domain experts while maintaining data quality (Please refer A.4 for more details on translation framework and expert validation).
+Additionally, the inclusion of text-based clinical data from BiMediXv1 Pieri et al.
+(2024) ensures the dataset retains robust language understanding capabilities while expanding its multimodal medical proficiency.
+This extensive dataset forms the foundation for advanced medical image-text alignment and conversational multimodal applications.
+For dataset composition, see Section A.7.
+Model Conversation Description CXR MRI Histology Gross CT Overall
+BiomedGPT 15.3 13.3 16.4 13.0 14.1 14.9 15.8 14.8
+MAIRA-2 19.1 27.0 45.6 13.7 13.8 15.0 19.1 21.2
+LLaVA-pp 34.3 36.6 44.7 33.3 34.7 30.2 31.5 34.9
+MiniGPT-Med 37.5 29.6 47.6 32.5 36.3 31.8 29.1 35.4
+LLaVA-Med 55.6 43.3 59.5 43.4 54.4 53.9 51.0 52.4
+Dragonfly-Med 59.2 34.2 67.0 51.2 53.7 42.6 48.3 52.7
+[…]
+Table 3: BiMed-MBench English Evaluation
+Model Conversation Description CXR MRI Histology Gross CT Overall
+BiomedGPT 11.1 11.2 11.4 10.8 11.5 11.3 11.1 11.2
+MAIRA-2 14.0 12.5 25.6 10.6 12.8 11.6 12.7 14.6
+MiniGPT-Med 21.6 12.6 23.7 12.7 32.0 15.8 14.9 20.2
+LLaVA-Med 23.9 29.4 31.2 25.3 24.8 23.4 26.4 26.2
+LLaVA-pp 29.0 27.8 33.2 25.0 33.0 25.8 25.8 28.7
+Dragonfly-Med 32.8 19.9 31.9 25.7 33.0 24.0 31.7 29.5
+[…]
+Table 4: BiMed-MBench Arabic Evaluation
+2.2 Medical Instruction Tuning
+To enable BiMediX2’s robust capabilities in both bilingual and multimodal medical tasks, we employ a two-stage training process that ensures precise alignment of visual and textual representations while adapting the language model for complex medical instruction tasks (see Fig 2).
+Stage 1: Medical Concept Alignment: In the first stage, we finetune the Projector alone to align visual embeddings to the language embedding space.
+The training utilizes a comprehensive dataset of 467k image-caption pairs sourced from the LLaVA-Med Li et al.
+(2023) pretraining dataset.
+These pairs span diverse medical imaging modalities and captions that describe clinically relevant features.
+Stage 2: Multimodal Medical Instruction Alignment: The second stage finetunes the LoRAHu et al.
+(2021) adapters within the language model, enhancing its ability to process and generate multimodal medical instructions.
+For this, we utilize our BiMed-V 1.6M bilingual multimodal instruction set, which comprises carefully crafted English and Arabic prompts paired with corresponding visual and textual responses.
+This dataset enables the model to learn nuanced instructions across a wide array of medical domains, from radiology to pathology, in a bilingual context.
+Through these two stages, BiMediX2 achieves seamless integration of bilingual and multimodal capabilities, enabling it to deliver accurate and context-aware medical insights in both English and Arabic, tailored to a variety of clinical scenarios.
+We have used LLaMA 3.1 (8B, 70B) and Phi-3.5 V Abdin et al.
+(2024) as base models to obtain BiMediX2 8B, BiMediX2 70B, and BiMediX2 4B variants, respectively.
+For further details on model and training configurations, see Section A.6.
+3 Experiments
+In the literature, evaluating medical language models predominantly involves multiple-choice question-answering tasks, with accuracy as the performance metric.
+We employed the EleutherAI Gao et al.
+(2024) evaluation framework for evaluating text-only medical tasks following Pieri et al.
+(2024).
+For our bilingual and multimodal evaluations, we utilized several benchmarks and tools.
+We leveraged our BiMed-MBench for assessing the correctness, helpfulness, relevance and accuracy of the model in a bilingual medical multimodal context.
+Additionally, we evaluated the model on VQA using datasets such as Path-VQA, SLAKE, and Rad-VQA.
+For report generation, we used the MIMIC-CXR Johnson et al.
+(2019) dataset, and for report summarization, we used the MIMIC-III Johnson et al.
+(2016) dataset.
+These evaluations were conducted using the MultiMedEval Royer et al.
+(2024) toolkit, ensuring a comprehensive assessment of the model’s multimodal capabilities.
+3.1 Medical LLM Benchmarks
+We assess our models on text-based clinical benchmarks using PubMedQA, MedMCQA, MedQA, USMLE, Medical MMLU, and UPHILL OpenQA. For further details, see Section A.3.
+Figure 3: Qualitative Examples of BiMediX2 for Medical Image Understanding in a Conversational Context.
+Figure 4: Performance comparison on UPHILL OpenQA Kaur et al. (2023), assessing the model’s ability to address false medical claims at different presupposition levels.
+3.2 Medical VLM Benchmarks
+BiMed-MBench: We introduced the first bilingual GPT-4o-based medical LMM benchmark, consisting of 386 medical queries spanning various medical imaging modalities.
+The test dataset includes conversational interactions and detailed descriptions for modalities such as chest X-rays (CXR), MRI, histology, gross and CT scans derived from LLaVA-Med Li et al.
+(2023).
 GPT-4o evaluates the correctness of model responses based on the provided image context and caption.
-The reference prediction, serving as the upper-bound answer for the teacher model, is generated using GPT-4 (Achiam et al.
-(2023)).
+The reference prediction, serving as the upper-bound answer for the teacher model, is generated using GPT-4 Achiam et al.
+(2023).
 For Arabic evaluations, the ground truth is derived from the upper-bound, translated by GPT-4o, and all test samples are comprehensively verified by medical experts for quality assurance.
@@
 These datasets include a mix of open-ended and close-ended questions, designed to test the model’s ability to interpret and respond to queries based on medical images.
-This evaluation ensures the model’s effectiveness in handling a variety of visual and textual inputs.
-Report Generation: The MIMIC-CXR (Johnson et al.
-(2019)) dataset, which includes de-identified radiology reports associated with chest X-rays, is utilized to assess the model’s performance in generating the findings section of medical reports.
-For this task, the input consists of one or more radiology images related to a case, followed by a prompt such as, “<<<image>>> <<<image>>> Please caption this scan with findings and impressions.” This setup evaluates the model’s ability to generate accurate and coherent medical reports based on visual inputs.
-Report Summarization: We conduct an evaluation of report summarization using the MIMIC-III (Johnson et al.
-(2016)) dataset.
-The task involves generating the impressions section of a radiology report based on the findings section.
-To achieve this, the model is provided with the free-text findings along with a task prompt, “Summarize the findings”.
-This assessment focuses on the model’s ability to distill comprehensive medical information into a concise and precise summary, which is essential for clear and effective medical communication.
-4 Results
-LLM Evaluation
-The performance of BiMediX2 and other models on various medical benchmarks is presented in Fig 3 and Tab 2.
-Our BiMediX2 70B achieved the highest average score of 84.6%, outperforming other models such as GPT-4 (82.9%) and Llama-3-Med42-70B (83.0%).
-BiMediX2 70B exhibited superior average performance across multiple datasets, including Medical MMLU, MedMCQA, MedQA, USMLE, and PubMedQA, demonstrating its strong understanding of medical contexts.
-Model Conversation Description CXR MRI Histology Gross CT Overall
-BiomedGPT 15.3 13.3 16.4 13.0 14.1 14.9 15.8 14.8
-LLaVA-pp 34.3 36.6 44.7 33.3 34.7 30.2 31.5 34.9
-MiniGPT-Med 37.5 29.6 47.6 32.5 36.3 31.8 29.1 35.4
-LLaVA-Med 55.6 43.3 59.5 43.4 54.4 53.9 51.0 52.4
-Dragonfly-Med 59.2 34.2 67.0 51.2 53.7 42.6 48.3 52.7
-BiMediX2 8B 64.9 54.5 71.7 56.8 62.5 61.4 58.9 62.2
-Table 3: BiMed-MBench English Evaluation
-Model Conversation Description CXR MRI Histology Gross CT Overall
-BiomedGPT 11.1 11.2 11.4 10.8 11.5 11.3 11.1 11.2
-MiniGPT-Med 21.6 12.6 23.7 12.7 32.0 15.8 14.9 20.2
-LLaVA-Med 23.9 29.4 31.2 25.3 24.8 23.4 26.4 26.2
-LLaVA-pp 29.0 27.8 33.2 25.0 33.0 25.8 25.8 28.7
-Dragonfly-Med 32.8 19.9 31.9 25.7 33.0 24.0 31.7 29.5
-BiMediX2 8B 54.3 36.2 61.4 44.6 51.5 43.5 50.8 50.5
-Table 4: BiMed-MBench Arabic Evaluation
-UPHILL OpenQA Evaluation
-Fig 4 illustrates the performance comparison on the UPHILL OpenQA benchmark.
-BiMediX2 70B achieved the highest overall factual accuracy of 60.6%, and the second highest being BiMediX2 8B (56.1%), surpassing other models such as GPT-4 (51.5%), Meditron 70B (49.6%), and Med42 (53.5%).
-This highlights BiMediX2’s effectiveness in discerning and correcting misinformation in medical contexts.
-Figure 4: Performance comparison on UPHILL OpenQA (Kaur et al. (2023)), assessing the model’s ability to address false medical claims at different presupposition levels.
-BiMed-MBench Evaluation
-Tables 3 and 4 present the evaluation results of BiMediX2 and other medical LMMs on the English and Arabic BiMed-MBench benchmark, respectively.
-BiMediX2 8B demonstrated superior performance in both English (overall score of 62.2%) and Arabic (overall score of 50.5%) evaluations, outperforming other models.
-Fig 1 compares the performance of state-of-the-art medical LMMs on our BiMed-MBench evaluation in a bilingual context.
-This indicates BiMediX2’s strong bilingual capabilities and its effectiveness in handling medical conversations and descriptions across different imaging modalities.
-Medical VQA Benchmark
-Tab 5 shows the performance of BiMediX2 and other models on the Medical VQA benchmark using the MultiMedEval toolkit.
-Our BiMediX2 8B achieves the highest average score of 0.611, outperforming other models across datasets such as Rad-VQA, Slake-VQA, and Path-VQA.
-This demonstrates BiMediX2’s proficiency in visual question answering, a critical task in medical diagnostics.
+The evaluation employs accuracy, precision, recall, and F1 score to measure the correctness and completeness of the model’s answers, which are critical for clinical reliability.
+Additionally, BLEU ensures the generated answers are naturally phrased and easy to understand for clinical communication.
 Dataset Metric RadFM LLaVA Med BioMedGPT MiniGPT-Med Phi-3.5 V BiMediX2 4B BiMediX2 8B
@@
 F1 ↑ 0.232 0.052 0.154 0.299 0.287 0.475 0.595
-Average 0.461 0.239 0.177 0.427 0.319 0.509 0.611
-Table 5: Medical VQA Benchmark (MultiMedEval Royer et al. (2024))
-Report Summarization
-Tab 6 presents the report summarization performance on the MIMIC-III dataset.
-BiMediX2 8B achieved the highest average score of 0.416, surpassing other models like LLaVA-Med (0.398) and Dragonfly-Med (0.110).
-This average score is derived as a unified metric by re-scaling BLUE-4* and RadCliQ* metrics.
-This indicates BiMediX2’s effectiveness in generating concise and accurate summaries of medical reports, a vital task for efficient healthcare communication.
-Report Generation
-Tab 7 shows the report generation performance on the MIMIC-CXR dataset.
-BiMediX2 8B achieved the highest average score of 0.235, outperforming other models like LLaVA-Med (0.192) and BioMedGPT (0.145).
-The average score is derived as a unified metric by re-scaling BLUE-4* and RadCliQ* metrics.
-This highlights BiMediX2’s capability in generating detailed and accurate medical reports from radiology images, a crucial task for diagnostic purposes.
-Dataset Metric LLaVA Med Dragonfly-Med BiMediX2 4B BiMediX2 8B
-MIMIC-III ROUGE-L↑ 0.185 0.072 0.209 0.205
-BLEU-1↑ 0.192 0.062 0.153 0.178
-BLEU-4↑* 0.520 0.000 0.410 0.449
-F1-RadGraph↑ 0.232 0.000 0.222 0.230
-RadCliQ↑* 0.753 0.247 0.923 0.918
-CheXbert vector↑ 0.600 0.326 0.633 0.593
-METEOR↑ 0.303 0.060 0.264 0.339
-Average 0.398 0.110 0.402 0.416
-Table 6: Report Summarization (MultiMedEval Royer et al. (2024))
-Dataset Metric RadFM LLaVA Med BioMedGPT BiMediX2 4B BiMediX2 8B
-MIMIC-CXR Report Generation F1-RadGraph↑ 0.042 0.048 0.000 0.083 0.098
-BLEU-1↑ 0.006 0.163 0.003 0.046 0.155
-BLEU-4↑* 0.000 0.060 0.000 0.042 0.016
-ROUGE-L↑ 0.065 0.125 0.012 0.131 0.153
-RadCliQ↑* 0.655 0.660 0.827 0.865 0.860
-CheXbert vector↑ 0.197 0.150 0.153 0.205 0.189
-METEOR↑ 0.053 0.137 0.016 0.107 0.174
-Average 0.145 0.192 0.145 0.211 0.235
-Table 7: Report Generation (MultiMedEval Royer et al. (2024))
-Qualitative Examples
-Medical Image Understanding in a Conversational Context: Fig 5 illustrates the capabilities of the BiMediX2 framework in analyzing medical images and providing detailed, context-aware responses.
-The top section showcases a conversation involving a sagittal CT scan of the lumbar spine.
-The model identifies the scan type and explains that it is a vertical slice of the body focusing on the lumbar spine.
-When asked about abnormalities, the model accurately identifies a fracture in the L4 vertebra, providing an explanation of the condition and its potential causes, such as trauma or stress.
-The bottom section features a color Doppler ultrasound scan of the left ovary, identifying a potential abnormality a left ovarian cyst with a solid component.
-Our model explains the imaging technique, names the organ, and discusses the detected abnormality, emphasizing the need for further evaluation.
-These examples demonstrate BiMediX2’s ability to interpret complex medical images and provide informative responses to assist in clinical decision-making.
-Figure 5: Qualitative Examples of our BiMediX2 for Medical Image Understanding in a Conversational Context.
+[…]
+Table 5: Medical VQA Benchmark MultiMedEval Royer et al. (2024)
+Report Generation: The MIMIC-CXR Johnson et al.
+(2019) dataset, which includes de-identified radiology reports associated with chest X-rays, is utilized to assess the model’s performance in generating the findings section of medical reports.
+For this task, the input consists of one or more radiology images related to a case, followed by a prompt such as, “<<image>> <<image>> Please caption this scan with findings and impressions." This setup evaluates the model’s ability to generate accurate and coherent medical reports based on visual inputs.
+Report Summarization: We conduct an evaluation of report summarization using the MIMIC-III Johnson et al.
+(2016) dataset.
+The task involves generating the impressions section of a radiology report based on the findings section.
+To achieve this, the model is provided with the free-text findings along with a task prompt, “Summarize the findings".
+This assessment focuses on the model’s ability to distill comprehensive medical information into a concise and precise summary, which is essential for clear and effective medical communication.
+In both report generation and summarization tasks, relying solely on lexical metrics such as BLEU, ROUGE-L, and METEOR is insufficient, as these do not guarantee clinical accuracy.
+Therefore, we complement them with clinically-informed metrics such as F1-RadGraph, CheXbert similarity, and RadCliQ.
+These metrics assess the preservation of key medical entities, relations, and overall clinical correctness, ensuring that the generated or summarized content aligns not just in phrasing but also in diagnostic intent and clinical relevance.
+Dataset Metric LLaVA Med Dragonfly-Med BiMediX2 4B BiMediX2 8B Med-PaLM M
+MIMIC-III ROUGE-L↑ 0.185 0.072 0.209 0.205 0.320
+BLEU-1↑ 0.192 0.062 0.153 0.178 0.154
+BLEU-4↑* 0.520 0.000 0.410 0.449 -
+F1-RadGraph↑ 0.232 0.000 0.222 0.230 0.347
+RadCliQ↑* 0.753 0.247 0.923 0.918 -
+CheXbert vector↑ 0.600 0.326 0.633 0.593 -
+METEOR↑ 0.303 0.060 0.264 0.339 -
+[…]
+Table 6: Report Summarization MultiMedEval Royer et al. (2024)
+(i) Baseline: VLM trained on non-medical Data.
+(ii) Stage-1: Aligning Medical images to the LLM input token space by learning an alignment layer and freezing the Language model and the Vision encoder using our pre-training dataset of 467k instructions.
+(iii) Stage-2: LoRA finetuning with our BiMed-V1.6M dataset comprising a text and image+text dataset across medical image modalities.
+Here the Baseline model is obtained by pretraining the projector on LCS-558K dataset following the LLaVA-pp Rasheed et al.
+(2024) repository.
+The baseline model lacks understanding of medical images which is introduced in our Stage-1 alignment training.
+Finally we LoRA finetune model to obtain our medically instruction tuned model.
+As observed in Tab 9, the text based medical evaluation scores do not change with Stage 1 training as we are only training the projector here and the language model performance is therefore consistent.
+Evaluation of English Model with Arabic Translation Pipeline: To further validate the necessity of bilingual instruction tuning, we employed a cascaded translation pipeline for evaluating our English instruction-tuned model (BiMediX2-8B ENG) on the BiMed-MBench Arabic benchmark using popular translation services such as from Google, Alibaba, and Bing.
+While this method yielded slight improvements over the base English model as shown in Tab 8, it consistently fell short compared to our bilingual model.
+These results demonstrate that translation pipelines alone are inadequate for generating medically accurate Arabic responses, highlighting the importance of dedicated bilingual training and instruction tuning.
+Further experiments are detailed in Section A.5.
+Model BiMed-MBench (Ara)
+BiMediX2-8B ENG 44.61
+BiMediX2-8B ENG
++ Translation (Google)
+44.92
+BiMediX2-8B ENG
++ Translation (Bing)
+45.05
+BiMediX2-8B ENG
++ Translation (Alibaba)
+45.33
+BiMediX2-8B BI 50.47
+Table 8: Comparison of Translation frameworks
+6 Qualitative Examples
+Medical Image Understanding in a Conversational Context: Fig 3 illustrates the capabilities of the BiMediX2 framework in analyzing medical images and providing detailed, context-aware responses in both English and Arabic.
+The top section highlights BiMediX2 analyzing a sagittal CT scan of the lumbar spine, correctly identifying the scan type and diagnosing an L4 vertebral fracture.
+It explains potential causes like trauma or stress and discusses clinical implications.
+In the bottom section, BiMediX2 accurately identifies a female reproductive organ, explains the imaging technique, and detects abnormalities, demonstrating bilingual capabilities in English and Arabic.
+Additional qualitative examples are provided in Section A.1, while Section A.2 presents its limitations.
+7 Conclusion
+BiMediX2 represents a leap forward in bilingual, multimodal medical AI, addressing the global need for accessible and inclusive healthcare solutions in both English and Arabic.
+By integrating text and visual modalities within a unified architecture, it enables seamless multi-turn interactions for diverse medical tasks, including medical image analysis and complex medical conversations.
+Key contributions include the comprehensive bilingual dataset, BiMed-V, which provides diverse multimodal medical instructions tailored for both languages, and the introduction of BiMed-MBench, the first bilingual GPT-4o-based medical benchmark, which showcases the model’s ability to excel in a wide range of expert-verified medical scenarios.
+BiMediX2 paves the way for inclusive, multilingual, and multimodal healthcare applications, significantly enhancing the accessibility and quality of medical assistance worldwide.
+Evaluation Baseline Stage-1 Stage-2
+BiMed-MBench [3] 34.9 47.6 62.2
+Clinical LLM Eval [2] 67.6 67.6 70.4
+Table 9: Multi-stage Training Ablation
+8 Limitations
+Despite its overall improvement, BiMediX2, like other generative language models, may experience issues such as hallucinations, toxicity, and stereotypes.
+These issues stem from both the inherited limitations of the base models and the nature of the pretraining data.
+While we have conducted automatic and qualitative evaluations with medical professionals, we acknowledge that our model’s medical diagnoses and recommendations may not always be accurate.
+Extensive human evaluation is more reliable but expensive and time-consuming.
+The exploration of alternative solutions remains an important focus for ongoing research.
+Currently, our models lack explicit mechanisms to curb undesirable behaviors.
+Future work will focus on enhancing alignment and safety strategies to reduce risks associated with clinical deployment.
+On a brighter note, we believe that releasing our model weights could contribute to investigating and mitigating these risks through broader community engagement.
+In addition, our current model does not explicitly mitigate biases related to gender, ethnicity, or socioeconomic status in medical contexts.
+We recognize the critical importance of fairness in AI-driven healthcare solutions, and acknowledge that such biases could perpetuate or exacerbate disparities in medical decision-making.
+As part of future work, we plan to conduct targeted bias assessments and explore debiasing strategies.
+We also believe that open-sourcing our model will facilitate further research into bias detection and mitigation.
+Another current limitation is that our model is trained exclusively on Modern Standard Arabic (MSA).
+This restricts its applicability in real-world clinical settings where dialectal Arabic, such as Emirati, Saudi, Egyptian, or Moroccan is more commonly used.
+In future iterations of this work, we aim to expand our training corpus to include a diverse range of Arabic dialects, thereby enhancing the model’s utility and inclusivity across Arabic-speaking populations.
+9 Safety and Ethical Implications
+We recognize the significant societal impact of BiMediX2 and emphasize the importance of ethical considerations and transparency.
+This work is intended for research purposes only and is not ready for clinical or commercial use.
+Ensuring the model’s accuracy and reliability is crucial, as incorrect medical advice could have serious health consequences.
+Robust validation and quality control measures are essential to minimize errors.
+Ethical considerations include protecting patient privacy and ensuring the confidentiality of medical data.
+The model must comply with relevant data protection regulations and ethical guidelines.
+Addressing potential biases in the model’s outputs is also critical for ensuring fair and equitable healthcare outcomes.
+Collaboration with patients, medical professionals, and ethicists is essential for ethical oversight and further research to ensure safety and accuracy in clinical settings.
+By acknowledging and addressing these considerations, we can continue to refine BiMediX2 for safe and effective use in healthcare.
+10 Acknowledgement
+We would like to thank Dr. Omair Mohammed, Dr. Mohammed Zidan, Dr. Vishal Thomas Oommen and Dr. Mehrubin Kurpat for their contribution in verification of medical responses.
+The computations were enabled by resources provided by LUMI hosted by CSC (Finland) and LUMI consortium, and by Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the NSC.
+This work is partially supported by the Meta Llama Impact Innovation Award, the Meta Regional Research Grant, the Google Research Award, the NVIDIA Academic Grant, and the MBZUAI-WIS Research Grant (P008).
+These recognitions highlight our commitment to advancing AI-driven healthcare solutions.
+Appendix A Appendix
+A.1 Additional Qualitative Examples
+Medical Image Understanding in a Conversational Context: Fig 8 illustrates the capabilities of the BiMediX2 framework in analyzing medical images and providing detailed, context-aware responses in both English and Arabic.
+The top section showcases a conversation involving a 3D CT scan of the chest.
+The model identifies the scan type and explains that it uses X-ray technology to create detailed cross-sectional images, which are then reconstructed into 3D images.
+When asked about abnormalities, the model accurately identifies multiple rib fractures.
+It further clarifies that these fractures are present on both the left and right sides of the chest.
+And in the bottom section, BiMediX2 accurately identifies the organ, explains the imaging technique, and detects abnormalities, providing valuable insights that can aid in the diagnosis and treatment of conditions related to the female reproductive system.
+This particular example showcases BiMediX2’s capability to converse in both English and Arabic simultaneously, depending on the input query.
 Medical Image Understanding in a Conversational and Bilingual (Arabic) Setting: Fig 10 illustrates BiMediX2’s ability to understand medical imagery and converse in Arabic, showcasing its bilingual capabilities.
@@
 This particular example showcases BiMediX2’s capability to converse in both English and Arabic simultaneously.
+Medical Image Understanding of our BiMediX2 in Multidomain: Fig 11 showcases BiMediX2’s versatility and accuracy across various medical imaging modalities.
+In the top section, it identifies key structures in a scanning electron micrograph of a mosquito head.
+In subsequent sections, it correctly interprets an MRI of a parotid tumor, a histology slide of adipose tissue (including stain type), a chest X-ray with pneumothorax, and a CT scan of the abdomen, pinpointing adrenal abnormalities.
+These examples demonstrate BiMediX2’s capability to analyze and diagnose diverse imaging types, making it a valuable tool across medical specialties.
+Figure 5: Data Translation Framework
+A.2 Model Failure Cases
+We present a common failure case observed in both English and Arabic settings. The example is shown in Fig 7.
+In the English-language evaluation, a common source of error was the model’s difficulty in distinguishing between visually similar but clinically distinct fracture types.
+As shown in Fig 7, while the model response is fluent and medically coherent, it contains a key factual error: the misclassification of an avulsion fracture as a burst fracture.
+Notably, this response was initially accepted by several medical professionals due to the similarity in language and presentation.
+However, a trained radiologist later identified the subtle radiographic distinctions, confirming the ground truth diagnosis as an avulsion fracture.
+This highlights the model’s challenge in handling nuanced diagnostic distinctions that require expert-level domain knowledge.
+In the Arabic benchmark, one of the most frequent error modes involved incorrect localization of the affected area as illustrated in Fig 7.
+The response demonstrates a localization mismatch where the model incorrectly identifies C2–C3 as the affected vertebrae instead of the correct C6–C7 levels.
+Such errors, especially in high-stakes medical applications, can have serious clinical implications, reinforcing the importance of precise anatomical grounding in medical AI systems.
+A.3 Clinical LLM Benchmarks
+PubMedQA Jin et al.
+(2019) is a question-answering dataset derived from biomedical research papers on PubMed.
+The task involves answering ’yes’, ’no’, or ’maybe’ based on question sourced from the title of a research paper and a context from the abstract.
+Our analysis focuses on the PQA-L subset, which includes 500 manually annotated QA pairs requiring in-depth biomedical reasoning.
+MedMCQA Pal et al. (2022) consists of 4,183 multiple-choice questions from Indian AIIMS and NEET PG medical entrance exams, assessing professional medical knowledge and language comprehension.
+MedQA Jin et al.
+(2021) features multiple-choice questions from medical board exams in the US, Mainland China, and Taiwan.
+Our study focuses on the USMLE portion (1,273 test samples), requiring multi-step reasoning and evidence retrieval.
+USMLE Han et al.
+(2023) is a self-assessment test from the United States Medical Licensing Examination (Step 1, Step 2, and Step 3).
+We use the MedAlpaca version, which excludes image-based questions and contains 325 test samples.
+Medical MMLU Hendrycks et al.
+(2020) is a collection of six datasets covering 1,089 test questions across Clinical Knowledge, College Biology, College Medicine, Medical Genetics, Professional Medicine, and Anatomy.
+UPHILL OpenQA Kaur et al.
+(2023) evaluates LLMs’ accuracy in handling health-related queries with varying presuppositions.
+Our analysis focuses on zero-shot models’ ability to refute false health claims, a critical factor in combating misinformation.
+In this context, the accuracy refers to the model’s effectiveness in accurately refuting false health-related claims at different presupposition levels.
+Fig  6 compares state-of-the-art medical LLMs and LMMs on clinical benchmarks.
+BiMediX2 70B achieved the highest average score (84.6%), surpassing GPT-4 (82.9%) and Llama-3-Med42-70B (83.0%), demonstrating its superior performance across diverse medical tasks.
+The stacked bars illustrate performance across individual datasets, highlighting BiMediX2’s strong understanding of medical contexts.
+A.4 Data Translation Framework
+To construct a high-quality bilingual dataset, we develop a robust data translation framework that leverages GPT-4o for translating English medical content into Arabic, followed by expert human verification to ensure contextual and terminological accuracy.
+This approach supports effective training and benchmarking in a bilingual context.
+We first translate our English multimodal instruction set to Arabic using GPT-4o.
+A random subset of this translated data is passed to Bilingual medical doctors to verify the quality of the translation.
+Doctors are asked to report the number of samples requiring minor (e.g rewording, formatting, or stylistic edits) / major (e.g incorrect clinical terms or misinterpretations) revision in the Arabic translation.
+They are required to look for the overall consistency of the translated text and the accuracy of medical terminologies.
+For our BiMed-MBench bilingual benchmark doctors were asked to manually correct these inaccuracies in the translated text.
+On average 22% of the samples required minor corrections and re-formatting, while only 5% of the samples required major corrections in medical terms.
+The verification process involved 10 medical experts from three different countries, ensuring representation from both native Arabic and English speakers.
+These experts specialized in various medical domains, including radiology (MD Radiology), endocrinology, neurosurgery, general practice, histopathology (MS Pathology), and general medicine.
+To ensure rigorous validation, a multi-reviewer setup was adopted for a randomly selected subset of the benchmark.
+Each sample was independently reviewed by multiple doctors, which surfaced occasional inconsistencies, particularly in ambiguous or borderline cases.
+To resolve such discrepancies, we introduced a structured adjudication protocol.
+Conflicting evaluations were circulated among the experts, followed by focused discussion sessions in which reviewers examined each other’s rationale.
+This collaborative refinement process significantly reduced inconsistencies and improved the reliability of the benchmark.
+A.5 Additional Experiments
+Model BiMed-MBench (Ara)
+BiMediX2-8B ENG 44.6
+BiMediX2-8B ARA 46.3
+BiMediX2-8B BI 50.5
+Table 10: Bilingual vs Monolingual model training
+Hyperparameter Stage 1 Stage 2
+Number of Epochs 1 1
+Train Batch Size (per device) 1 4
+Gradient Accumulation Steps 1 4
+Learning Rate 1.0×10-3 2.0×10-4
+Optimizer Adam Adam
+Weight Decay 0 0
+Warmup Ratio 0.03 0.03
+Learning Rate Scheduler Cosine Cosine
+Precision bfloat16 bfloat16
+PEFT None LoRA
+LoRA Rank - 8
+LoRA Alpha - 16
+Multi-Modal Projector LR - 2.0×10-5
+Table 11: Model training Hyperparameters
+A.5.1 Human Expert Evaluation
+To assess the clinical quality of model outputs, we conducted a human expert evaluation using a blind review setup.
+For a randomly selected subset of questions from the BiMed-MBench benchmark, responses were generated by three models: BiMediX2, Dragonfly-Med, and LLaVA-Med.
+The outputs were anonymized and labeled as Model A, Model B, and Model C, with no identifiers provided to the reviewers.
+Medical experts were asked to evaluate the responses against the provided ground truth descriptions for each question.
+The evaluation focused on determining which model produced the most accurate, clinically relevant, and clear explanation of the medical image.
+The results demonstrate a strong preference for BiMediX2, which was selected as the best response in 76.9% of the cases.
+In comparison, Dragonfly-Med was preferred in 15.4% of the cases, and LLaVA-Med in 7.7%.
+Evaluation Protocol for Medical Experts: Your task is to evaluate the responses provided by three AI models based on a given medical image description (Ground Truth).
+Follow these steps to make your selection:
+1) Read the Ground Truth: Carefully review the provided description of the medical image. This serves as the reference for an accurate and detailed response.
+2) Assess the Model Responses: Examine the three model-generated responses (Model A, Model B, and Model C).
+Compare their content with the Ground Truth, focusing on the accuracy, completeness, and relevance of the description.
+3) Select the Best Response: Choose the model response that best aligns with the Ground Truth in terms of:
+- Clinical Accuracy: Does the response correctly describe the key findings in the image?
+- Relevance: Does the response stay focused on the specific details highlighted in the Ground Truth?
+- Clarity: Is the explanation well-structured and easy to understand
+4) Submit Your Choice: After evaluating the responses, select the one that provides the most accurate and comprehensive explanation.
+Dataset No. of samples
+PubMedQA 210169
+MedMCQA 182712
+MedQA 20691
+Single Turn QA 119879
+Multi-Turn Conversation 133134
+PMC-MCQ 80000
+LLaVA-Med-Subset-to-Conv 11616
+Rad-VQA 1796
+Slake-VQA 9835
+PMC-VQA 80000
+Path-VQA 19654
+LLaVA-Med-to-QA 163463
+MedQA (Ara) 11210
+PubMedQA (Ara) 115773
+MedMCQA (Ara) 97523
+Single Turn QA (Ara) 156254
+Multi-Turn Conversation (Ara) 64235
+LLaVA-Med-to-QA (Ara) 163463
+PMC-MCQ (Ara) 50000
+[…]
+Table 12: BiMed-V-1.6M Dataset Composition
+A.5.2 Comparison of Bilingual vs Monolingual Training
+The results in Tab 10 highlight the impact of language setting on model performance for Arabic medical evaluation tasks.
+The bilingual model (BiMediX2-8B BI), trained on both English and Arabic instruction data, achieves the highest score of 50.5, outperforming both the Arabic-only model (46.3) and the English-only model (44.6) on the Arabic subset of BiMed-MBench.
+This demonstrates that bilingual instruction tuning provides complementary knowledge transfer, enabling the model to better generalize in Arabic.
+The performance gain over the Arabic-only model further emphasizes the value of incorporating English medical knowledge during training, even when the target evaluation is in Arabic.
+A.6 Model Configurations
+All traning experiments were conducted with 8× AMD Instinct MI200 GPUs (each with 64 GB of VRAM).
+The training process is structured in two stages: Stage 1 training requires approximately 20 hours, and Stage 2 training requires approximately 32 hours on this setup.
+The complete set of training hyperparameters are provided in Tab 11.
+We also compare the trade-offs between our model sizes along with their corresponding performance and medical accuracy in Tab 13.
+We compare the Token Throughput measured in (Tokens/second), Mean Time to First Token (TTFT) (ms), GPU VRAM consumed and Medical Accuracy based on our evaluation benchmarks.
+These benchmarks are run using the vLLM benchmarking suite Kwon et al.
+(2023) on 4xNVIDIA RTX A6000 (48GB) GPUs.
+A.7 BiMedV-1.6M Dataset
+The BiMedV-1.6M dataset comprises 1,691,407 samples across English and Arabic, covering text and image-text QA tasks, supporting multilingual and multimodal medical research.
+The dataset composition is presented in Tab 12, Fig 9.
+Model Token Throughput (tok/s) Mean TTFT (ms) GPU VRAM (GB) Medical Accuracy
+BiMediX2-4B 263.05 180.5 42 50.5
+BiMediX2-8B 187.14 241.9 44 70.4
+BiMediX2-70B 67.61 863.6 170 84.6
+Table 13: Performance comparison of BiMediX2 variants
+Figure 6: State of the art comparison of models in Clinical LLM Benchmarks
+Figure 7: Bilingual failure case: Model labels avulsion fracture as burst (English), and misidentifies C2–C3 instead of C6–C7 (Arabic).
+Figure 8: Qualitative Examples of BiMediX2 for Medical Image Understanding in a Conversational Context.
+Figure 9: BiMed-V-1.6M Dataset Distribution.
+The sunburst chart illustrates the hierarchical composition of the dataset, categorizing samples by language (English and Arabic), modality (text and image-text), and task type (QA and VQA).
+The inner rings represent language and modality divisions, while the outer rings detail specific datasets included in BiMedV-1.6M.
 Figure 10: Qualitative Examples of our BiMediX2 for Medical Image Understanding in a Conversational and Bilingual (Arabic) Setting: The top section translates to; User: What is the organ shown in the image?
@@
 BiMediX2: The abnormality in this ultrasound image is the presence of a multilocular ovarian cyst.
-Medical Image Understanding of our BiMediX2 in Multidomain: Fig 7 showcases BiMediX2’s versatility and accuracy across various medical imaging modalities.
-In the top section, it identifies key structures in a scanning electron micrograph of a mosquito head.
-In subsequent sections, it correctly interprets an MRI of a parotid tumor, a histology slide of adipose tissue (including stain type), a chest X-ray with pneumothorax, and a CT scan of the abdomen, pinpointing adrenal abnormalities.
-These examples demonstrate BiMediX2’s capability to analyze and diagnose diverse imaging types, making it a valuable tool across medical specialties.
 Figure 11: Qualitative Examples for Medical Image Understanding of our BiMediX2 in Multidomain: Capability of BiMediX2 to understand and interpret various medical imaging modalities and provide accurate responses.
@@
 These examples highlight BiMediX2’s versatility and effectiveness in medical image analysis and diagnosis.
-5 Conclusion
-BiMediX2 represents a leap forward in bilingual, multimodal medical AI, addressing the global need for accessible and inclusive healthcare solutions in both English and Arabic.
-By integrating text and visual modalities within a unified architecture, it enables seamless multi-turn interactions for diverse medical tasks, including medical image analysis and complex medical conversations.
-Key contributions include the comprehensive bilingual dataset, BiMed-V, which provides diverse multimodal medical instructions tailored for both languages, and the introduction of BiMed-MBench, the first bilingual GPT-4o-based medical benchmark, which showcases the model’s ability to excel in a wide range of expert-verified medical scenarios.
-BiMediX2 paves the way for inclusive, multilingual, and multimodal healthcare applications, significantly enhancing the accessibility and quality of medical assistance worldwide.
-6 Limitations
-Despite its overall improvement, BiMediX2, like other generative language models, may experience issues such as hallucinations, toxicity, and stereotypes.
-These issues stem from both the inherited limitations of the base models and the nature of the pretraining data.
-While we have conducted automatic and qualitative evaluations with medical professionals, we acknowledge that our model’s medical diagnoses and recommendations may not always be accurate.
-Extensive human evaluation is more reliable but expensive and time-consuming.
-The exploration of alternative solutions remains an important focus for ongoing research.
-Currently, our models lack explicit mechanisms to curb undesirable behaviors.
-Future research will focus on enhancing security and alignment strategies.
-On a brighter note, we believe that releasing our model weights could contribute to investigating and mitigating risks tied to application uses.
-7 Safety and Ethical Implications
-We recognize the significant societal impact of BiMediX2 and emphasize the importance of ethical considerations and transparency.
-This release is intended for research purposes only and is not ready for clinical or commercial use.
-Ensuring the model’s accuracy and reliability is crucial, as incorrect medical advice could have serious health consequences.
-Robust validation and quality control measures are essential to minimize errors.
-Ethical considerations include protecting patient privacy and ensuring the confidentiality of medical data.
-The model must comply with relevant data protection regulations and ethical guidelines.
-Addressing potential biases in the model’s outputs is also critical for ensuring fair and equitable healthcare outcomes.
-Collaboration with patients, medical professionals, and ethicists is essential for ethical oversight and further research to ensure safety and accuracy in clinical settings.
-By acknowledging and addressing these considerations, we can continue to refine BiMediX2 for safe and effective use in healthcare.
-8 Acknowledgement
-We would like to thank Dr. Omair Mohammed, Dr. Mohammed Zidan and Dr. Vishal Thomas Oommen for their contribution in verification of medical responses.
-The computations were enabled by resources provided by LUMI hosted by CSC (Finland) and LUMI consortium, and by Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the NSC.
-We are grateful to the Meta Llama Impact Innovation Awards for recognizing BiMediX2, as one of the winners awarded in October 2024.
-This recognition highlights our commitment to advancing AI-driven healthcare solutions.